OpenStudy (anonymous):

Please solve each of the inequalities below and describe the graph. 1. x – 3 > –9 2. –2x + 1 ≤ –11 3. 10 < –3x + 1 4. 2(x + 5) > 8x – 8 5. –2(x – 3) ≥ 5 – (x + 3)

OpenStudy (anonymous):

1. x - 3 > -9 x > -9 + 3 x > - 6 open circle on -6, shading to the right. 2. -2x + 1 <= - 11 -2x <= -11 - 1 -2x <= - 12 x => -12/-2 (change signs when dividing by a negative) x = > 6 closed circle on 6, shading to the right 3. 10 < -3x + 1 10 - 1 < - 3x 9 < - 3x 9/-3 > x -3 > x open circle on -3, shading to the left. 4. 2(x + 5) > 8x - 8 2x + 10 > 8x - 8 2x - 8x > -8 - 10 -6x > - 18 x < -18/-6 x < 3 open circle on 3, shading to the left. 5. -2(x - 3) >= 5 - (x + 3) -2x + 6 > = 5 -x - 3 -2x + x > = 5 - 3 - 6 -x > = -4 x < = 4 closed circle on 4, shading to the left. Just remember, if there is an equal sign in the problem, it will always be a closed circle. Also, if it is less then, shading goes to the left. If it is greater then, shading goes to the right.

OpenStudy (anonymous):

thanks so very much

OpenStudy (anonymous):

no problem....glad I could help :)